OpenStudy (anonymous):

divide the polynomials. -a^3b^2 + a^2b^2 - ab^3 over -a^2b^2

OpenStudy (lalaly):

\[\large{\frac{-a^3b^2+a^2b^2-ab^3}{-a^2b^2}} = \frac{-a^3b^2}{-a^2b^2}+\frac{a^2b^2}{-a^2b^2}-\frac{a^3b}{-a^2b^2}\]\[=\large{a-1+\frac{a}{b}}\]

OpenStudy (anonymous):

thank you again. i have a lot to do so i hope you like these.